Keck instruments
Download
1 / 53

Keck Instruments - PowerPoint PPT Presentation


  • 74 Views
  • Uploaded on

Keck Instruments. The WMKO Support Astronomer Group. General. New SA: Scott Dahm from Caltech FY08 upgrades slowed by unfilled 8 th SA position. Instrument teams continue to provide excellent support. Web-based documentation: Observer manuals; throughputs or S/N calculators

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' Keck Instruments' - sanam


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
Keck instruments

Keck Instruments

The WMKO Support Astronomer Group


General
General

  • New SA: Scott Dahm from Caltech

  • FY08 upgrades slowed by unfilled 8th SA position.

  • Instrument teams continue to provide excellent support.

  • Web-based documentation:Observer manuals; throughputs or S/N calculators

  • Electronic Observers’ NewsletterPositive feedback


The observing experience
The Observing Experience

  • New “desktop”: FVWM

    • OpenLook no longer supported by Sun

    • ssh on all computers

    • VNC integration

  • RemOps I noise reduction

  • Simultaneous DVD writing

  • Better chairs

  • Scanners available.



Nirspec

NIRSPEC

Jim Lyke


Nirspec1
NIRSPEC

  • From last year’s presentation:

    • Rotator-echelle interaction (no action)

    • Faulty SLIT init switch (now working)

    • Server crashes (fiber issues)

    • Bad SCAM quadrant

    • Warm & pump instead of window replacement

    • MAGIQ

  • New this year:

    • Noise investigation (Tom Greene visit)

    • Documentation update team identified

      • News is up-to-date


Nirspec magiq

Camera head

Fold mirror

Focus stage

Dewar window

Filter wheels

NIRSPEC MAGIQ

  • MAGIQ camera installed and running every night

    • IQM not implemented (eng time on May 17)

    • Performance good, numbers after eng time

    • Some software & procedural issues


Nirspec bad scam
NIRSPEC Bad SCAM

  • Bad SCAM quadrant

    • Reads all 0s

    • Affects all slits

  • User-defined nods

  • Do we try to fix?

    • Warm & pump fall 2008

  • Similar problem prior to delivery

    • Nothing changed

    • “Fixed” by thermal cycle


Nirspec noise
NIRSPEC Noise

Plot of single rows

  • Noise Investigation

    • Tom Greene visit Apr 2008

    • Digital (white plot)

      • Replaced bad A/D

    • Analog (red plot)

      • Left-side, 1 in 8 rows

      • Cooling? Grounding?

  • No read noise change

    • CDS: 50-55 e-

    • MCDS (16): 12-15 e-


Nirspec ice

Window replacement + 3 months

Warm & pump + 12 months

(current)

3.45 mm

3.10 mm

2.65 mm

NIRSPEC Ice

Next warm & pump requested for fall 2008


Osiris

OSIRIS

Jim Lyke


Osiris1
OSIRIS

  • From last year’s presentation:

    • Efficiency (FOV & Acq.) – cookbook available

  • New this year:

    • March 2008 Service

    • Observers really like NIRC2/OSIRIS switching

    • Manual & web pages update


Osiris service

Stage limit

August 2007

OSIRIS Service

  • March 2008 Service

    • Add’l K-band filters w/ pupils for 100 mas

    • Removed redundant z-band filters

    • Fixed handling cart issue

    • Eng time for pupil performance (June 5)

  • Software updates

    • GUIs, scripts, DRP (mostly for new filters)

    • Strong relationship w/ UCLA team


Osiris2
OSIRIS

  • DRP

    • Observers note slowness (45-60s per cube)

    • Users at Keck, UCLA, CIT, UCB

  • NIRC2/OSIRIS switching

    • Allows hi-res imaging/spectra in one night

    • Quick switch (no tertiary move)

  • Manual & web pages

    • New pupil performance data not available yet


Nirc2

NIRC2

Al Conrad


Outline
Outline

  • Differential Tracking with LGSAO.

  • More NIRC2 characterization on the web.

  • Proposed Service Mission.


Differential tracking with lgsao
Differential Tracking with LGSAO

  • LGS-AO now allows differential motion between the tip-tilt reference and the laser guide star.

  • This mode allows use of an off-axis tip-tilt star for a nonsidereal science object.

  • First demonstrated with NIRC2 in mid-February (~30 mas rms jitter)

  • Successfully used for science in late March

  • Will be offered as a shared risk mode on future LGSAO nights (NIRC2 and OSIRIS).

  • “Vertical angle” mode coming soon.


Nirc2 characterization on the web
NIRC2 Characterization on the Web

Astrometry page:

  • Recommended fields

  • PA accuracy

  • Dewarp tool.

  • Limiting rates and magnitudes for the 6 differential tracking modes.

  • A preliminary page for collecting JHK image quality comparisons from pre- and post-NGWFC.


  • Proposed service mission
    Proposed Service Mission

    • Status

      • Submitted as an FY09 project.

      • Not funded. Earliest date therefore Oct 2009.

      • AOWG comments will be solicited.

    • Summary of Proposed Changes

      • Filters

        • L-wide (to be acquired)

        • 8 others, including Y (in hand)

      • Polarimetry

        • Wollaston prism for NIRC2 filter wheel

        • Wave plate for the AO bench


    Service mission science drivers
    Service Mission Science Drivers

    • L filter will improve efficiency for both imaging and spectroscopy

    • Y-band filter will improve multi-color photometry studies of

      • low mass binaries

      • solar system objects.

    • Polarimetry

      • galactic center flares

      • structure and composition of circumstellar disks.


    LRIS

    Marc Kassis


    Performance and upgrades
    Performance and Upgrades

    • Overall, LRIS is performing well

    • Less than 4% of time lost to faults

    • Observer documentation updated end of last year

    • Software upgrades

      • Compass rose deployed for the Blue side

      • Focus procedures updated

        • Script configures LRIS, turns on appropriate lamps, and acquires images

        • Long slit focusing acquisition and analysis procedures available and documented


    Atmospheric dispersion corrector adc
    Atmospheric Dispersion Corrector (ADC)

    • Released for routine observing in 2007B

    • Highly reliable operation; no nighttime faults encountered

    • Observer documentation was developed and published on web

    • Investigating a problem that puts ADC prisms into secondary limit during installation

    • ADC does not limit the precision of polarization measurements taken with LRISp

      • Using normal observing techniques

      • Analysis available on the web


    Lris broken i band filter
    LRIS Broken I-Band Filter

    • Cancelled three LRIS nights

    • Metal to filter contact may result in chipping.

    • Filter chipped during 2006 earthquakes.

    • Crack propagated as filter rattled.

    • Shards jammed in grism mechanism


    Filter repair
    Filter Repair

    • Filter particles found

      • In grism carousel

      • On three grisms and cells

      • Main LRIS compartment

    • Vacuumed filter and grism carousels and main LRIS compartment

    • Repaired chipped filters

      • 50% of red and blue filters had small chips at edges.

      • Ground out chipped areas

      • Taped filter edges

    • Used “First Contact” to clean optical surfaces


    Replacement filter
    Replacement filter

    • September 2007–queried LRIS observers about I-band filter replacement

    • Most observers preferred an I filter similar to Pan-STARRS i-band.

    • Z-band desired by community

    • I-band ordered from Barr Associates and will be delivered September 2008


    Calibration lamps

    Hg, Zn, Cd lamps

    1200/3400 grism

    Tint=10s

    ADUs

    Angstroms

    Fe/Ar lamp

    1200/3400 grism

    Tint=30s

    ADUs

    Angstroms

    Calibration Lamps

    • Improve calibrations at 3000–4000 Angstroms

    • Install Fe/Ar and deuterium lamps for calibrating blue side

    • New lamp controller purchased

    • New lamp holder designed and being fabricated

    • Next step: test controller and new software in the lab


    Upgrade slitmask frames and pickoff mirrors
    Upgrade Slitmask Frames and Pickoff Mirrors

    • Goals

      • Improve mechanical reliability

      • Increase slitmask alignment efficiency

    • Slitmask frames redesigned

    • New bars to hold pickoff mirror designed

    • ME group will machine and assemble the frames and bars using Keck’s new CNC milling machine

    • Engineering with the prototype scheduled for May-June 2008


    Slitmask prototype
    SlitmaskPrototype

    • No screws

    • Pins in feet for reliable positioning

    • Slitmask curvature maintained by frame


    Deimos

    DEIMOS

    Greg Wirth


    Deimos performance notes
    DEIMOS Performance Notes

    • DEIMOS continues to perform well overall

    • Reflective coatings are in good condition

    • Faults averaged 3.9% over last year, relatively low

    • Grating system showing signs of wear

      • Clamp-up problems on slider 4

      • Possible connection to FCS tracking failures

    • Low-level bias variations affect 2 of 8 science CCDs since earthquake; impact low


    Deimos time loss

    Total lost time in last year was 21 hours

    Most grating loss was due to 1 major failure (6.5 hours)

    DEIMOS Time Loss

    May 2007 – April 2008 by source


    Deimos improvements
    DEIMOS Improvements

    • Installed science-grade imaging flat

    • Milled new reflective variable longslit mask (5 slits w/ widths 0.7–1.5 arcsec)

    • Released custom arc-line plotting software tool for distribution

    • Published procedure for generating slitmask finder charts


    Deimos future priorities
    DEIMOS Future Priorities

    • Restore normal operation of grating system

    • Create combined longslit+focusing mask (free up 1 slitmask slot)

    • Troubleshoot CCD noise problem

    • Purchase replacement V filter (maybe)


    Esi status report

    ESI Status Report

    Greg Wirth


    Esi servicing mission 2
    ESI Servicing Mission 2

    • Prompted by May 2007 breakdown in imaging flat stage

    • Scope:

      • Repair broken coupling on imaging flat stage

      • Remove coating of titanium dust from optics

      • Machine shutter blades to reduce friction, dust

      • Troubleshoot collimator mirror actuators

      • Remove reflective optics for recoating

    • Job completed in Jul/Aug 2007

    • Total ESI expenditures in FY07:

      • $67k from WMKO (incl. labor, materials, Lick contract)

      • $30k from UCO/Lick (donated labor and materials)

    • Instrument now working well again!


    Esi optics
    ESI Optics

    • Key optics stripped and recoated with “Holy Grail” Ag-on-Al process at UCO/Lick

    • New coating has excellent reflectance

    • Decent longevity expected

    • Collimator substrate in poor condition; explains poor appearance of coating


    Scattered light from esi collimator
    Scattered Light from ESI Collimator

    Scattering due to poor condition of mirror surface (not a coating problem)


    Esi status
    ESI Status

    • Recent usage shows small uptick:

      • 2007A: 14 nights scheduled

      • 2007B: 11 nights scheduled

      • 2008A: 16 nights scheduled

    • Recent developments:

      • Fixed broken air cylinder on shutter (other unit already replaced in servicing mission)

      • Recoated optics working well


    Esi possible future actions
    ESI Possible Future Actions

    • Complete commissioning of ESI Integral Field Unit (IFU)

      • Low demand for this mode

      • Requires 1-2 engineering nights

      • Request SSC guidance on the priority


    Hires

    HIRES

    Grant Hill


    Hires recent work

    The past year has been a busy one for HIRES:

    • New Th/Ar lamp (original #1 died).

      • Better Argon/Neon balance.

      • Does not illuminate slit quite as fully.

    • HIRES electronics vault removed and much larger laser

    • enclosure put in place.

    • Repairs on aging mechanisms:

    • Optics covers Echelle stage motor control

    • Guider cooling Shutter

    • Oxygen sensor Shutter controller

    • Pneumatic actuators

    • Corrected software bug in corrector cover status

    • For several months, air purge desiccant was saturating

    • ~ weekly.

    HIRES Recent Work


    Hires current issues
    HIRES Current Issues

    • Gradients (<1%) in CCD biases.

    • Wavelength drifts at 1-2 pixel level (sporadic).

    • Runaway software process.

    • Humidity levels still being monitored.

    • Focus slowly relaxing towards pre-earthquake values

    • Optical coatings need monitoring.

    • Documentation. Remaining big items in order of priority:

      • Exposure time calculator.

      • Observers’ manual.

      • Echelle format simulator.

      • Th/Ar atlas/map.

      • Calibration advice.

      • Observing/post-observing checklists.



    Instrument pm project

    • FY08 includes resources to establish an instrument PM program.

    • We are not where we had hoped to be at this point though:

      • Mainly due to observatory resource shortages beyond

      • control of IPM team.

      • Somewhat under-resourced for FY08 to begin with.

    • But still some good news:

      • Work resuming on IPM project.

      • First draft of software requirements document written.

      • Compilation of PM task database begun.

      • Some PM’s are being done.

      • We are better resourced in FY09 than in FY08.

    Instrument PM Project


    Mainland observing

    Mainland Observing program.

    Gregory D. Wirth


    Mainland observing recent developments

    Supported sites: program.

    Caltech*

    LBNL*

    UCLA*

    UCSC*

    UCSD*

    UCB* (new)

    UCSB (new)

    UCR (new)

    Future sites:

    UCI (summer 2008)

    UCD (fall 2008)

    Caltech 2nd room

    UH-Manoa (maybe)

    *Specified sites support mainland-only observing mode

    Mainland Observing: Recent Developments


    Mainland observing usage performance
    Mainland Observing: program.Usage & Performance

    • Capabilities enhanced

      • Second set of ISDN backup lines installed

      • Enables mainland-only mode on both telescopes

    • Response improved

      • Summit→Mainland link now 1GB/s (>6x faster)

      • HQ→Summit link now 1GB/s (>20x faster)

    • Reliability outstanding

      • Internet outages rare last year

      • No observing time loss

    • Popularity increasing

      • Now used on ~1/3 of nights


    Mainland observing upgrades
    Mainland Observing: program. Upgrades

    • Hardware: moving to dedicated VNC computers at summit

    • Software: being made more usable and maintainable


    KOA program.

    Hien Tran


    KOA program.

    • Release of HIRES single CCD data

      • 2001–2004 data release early May

      • 1996–2000 data release June 30

    • Extracted spectra

      • Mosaic data only

      • Pipeline use of MAKEE

      • “Browse” product only

      • Release in August


    ad